OpenStudy (anonymous):

How do I simplify: m+3/m+2 x m+2/m+1 x 5m/m+1?

OpenStudy (anonymous):

How do i answer this type of question?

OpenStudy (anonymous):

those x means *?

OpenStudy (turingtest):

\[{m+3 \over m+2}\times{m+2 \over m+1}\times{5m \over m+1}\]is the problem???

OpenStudy (anonymous):

yes.

OpenStudy (turingtest):

okay, then the (m+2)'s cancel giving\[{m+3 \over m+1}\times{5m \over m+1}={5m^2+15m \over m^2+2m+1}\]This seems simplified to me.

OpenStudy (anonymous):

the answer is: 6m+3 ----- m+1

OpenStudy (anonymous):

either your given answer is wrong, or the question is wrong

OpenStudy (anonymous):

okay, the question is: m+3 ---- m+2

OpenStudy (anonymous):

[5m ^{2}+15m \div(m+1)^{2}\]

OpenStudy (anonymous):

or u can open the formula

OpenStudy (anonymous):

the question is: m+3 m+2 5m ---- x ---- + ---- m+2 m+1 m+1

OpenStudy (anonymous):

do I multiply m+3 to m+2 and 5m?

OpenStudy (anonymous):

u can cancl m+2

OpenStudy (anonymous):

ohh, okay.

OpenStudy (turingtest):

ah, a plus sign changes everything! again (m+2)'s cancel giving\[{m+3 \over m+1}+{5m \over m+1}={6m+3 \over m+1}\] we can just add across as you see.

OpenStudy (anonymous):

yep, now it's right

OpenStudy (anonymous):

noo, the question is: m+3 m+2 5m ---- x ---- + ---- m+2 m+1 m+1

OpenStudy (anonymous):

TuringTest wrote the correct answer, he jsut cancelled both m+2 :)

OpenStudy (anonymous):

that doesn't change a thing

OpenStudy (anonymous):

ohh, okay

OpenStudy (anonymous):

ohhh,i see it! Thank you!!